# Attention Mechanisms Don’t Learn Additive Models: Rethinking Feature Importance for Transformers

### v1

@@ Abstract
We address the critical challenge of applying feature attribution methods to the transformer architecture, which dominates current applications in natural language processing and beyond. Traditional attribution methods to explainable AI (XAI) explicitly or implicitly rely on linear or additive surrogate models to quantify the impact of input features on a model’s output. In this work, we formally prove an alarming incompatibility: transformers are structurally incapable to align with popular surrogate models for feature attribution, undermining the grounding of these conventional explanation methodologies. To address this discrepancy, we introduce the Softmax-Linked Additive Log-Odds Model (SLALOM), a novel surrogate model specifically designed to align with the transformer framework. Unlike existing methods, SLALOM demonstrates the capacity to deliver a range of faithful and insightful explanations across both synthetic and real-world datasets. Showing that diverse explanations computed from SLALOM outperform common surrogate explanations on different tasks, we highlight the need for task-specific feature attributions rather than a one-size-fits-all approach.
@@ 1 Introduction
The transformer architecture [55] has been established as the status quo in modern natural language processing [15, 38, 39, 54]. However, the current and foreseeable adoption of large language models (LLMs) in critical domains such as the judicial system [11] and the medical domain [26] comes with an increased need for transparency and interpretability. Methods to enhance the interpretability of an artificial intelligence (AI) system are developed in the research area of Explainable AI (XAI, 2, 20, 36, 9). A recent meta-study [42] shows that XAI has the potential to increase users’ understanding of AI systems and their trust therein. Feature attribution methods that quantify the contribution of each input to a decision outcome are among the most popular explanation methods, and a variety of approaches have been suggested for the task of computing such attributions [27, 40, 51, 31, 14, 34].
@@
GAM, ensembles, boosting GAM: ✓non-linearities ✗no interactions Removal-based, e.g. Shapley Values, Local GAM approx.
@@
Figure 1: Transformers cannot be well explained through additive models. Left: We exemplarily show the log-odds for the outputs of a BERT model and a linear Naïve-Bayes model (“linear”) trained on the IMDB movie review dataset. We pass two sequences to the model independently and in concatenation. While for the linear model, the output of the concatenated sequence can be described by the sum, this is not the case for BERT. We show that this phenomenon is not due to a non-linearity in this particular model but stems from a general incapacity of transformers to represent additive functions. Right: To overcome this difficulty, we propose SLALOM, a novel surrogate model specifically designed to better approximate transformer models.
@@
However, it remains hard to formally define the contribution of an input feature for non-linear functions. Recent work [22] has shown that many common explanation methods do so by implicitly or explicitly performing a local approximation of the complex black-box function, denoted as $f$, using a simpler surrogate function $g$ from a predefined class $\mathcal{G}$. For instance, Local Interpretable Model-agnostic Explanations (LIME, 40) or input gradient explanations [5] use a linear surrogate model to approximate the black-box $f$. An important implication of this view is the recovery property. This property implies that if the black-box function is already within the function class $\mathcal{G}$, we can effectively reconstruct the original function $f$ from the explanations. As an example, suppose that the black box function we consider is of linear form, i.e., $f({\bm{x}})={\bm{w}}^{\top}{\bm{x}}$. In this case, a gradient explanation as well as continuous LIME (C-LIME, 3) will recover the original model’s parameters up to an offset [22, Theorem 1]. Shapley value explanations [31] possess a comparable relationship: It is known that they correspond to the feature contributions of Generalized Additive Models (GAM, 7).
The significance of recovery properties lies in their role when explanations are leveraged to gain insights into the underlying data. Particularly in applications like XAI for scientific discovery, preserving the path from the input data to the explanation through a learned model is crucial. These guarantees, however, hold practical value only when the black-box model can be approximated by the chosen surrogate function class $\mathcal{G}$ at least within some local region.
In this study, we demonstrate that the transformer architecture, the main building block of LLMs such as the GPT models [39], is inherently incapable of learning additive and linear models on the input tokens, both theoretically and empirically. We formally prove that encoder-only and decoder-only transformers structurally cannot represent additive models due to the attention mechanism’s softmax normalization, which necessarily introduces token dependencies over the entire sequence length. An example is illustrated in Figure 1 (left).
@@
Our findings imply that additive explanations cannot be faithful to transformers. This is a significant oversight in current XAI practices which undermines the explanations’ recovery properties. As our results suggest that the role of tokens cannot be described through a single score, we introduce the Softmax-Linked Additive Log-Odds Model (SLALOM, cf. Figure 1, right), which represents the role of each input token in two dimensions: The token value describes the independent effect of a token, whereas the token importance provides a token’s interaction weight when combined with other tokens in a sequence. In summary, our work offers the following contributions over the related literature:
@@
    We theoretically and empirically demonstrate that common transformer architectures fail to represent GAMs and linear models on the input tokens, reducing explanation faithfulness
@@
    To mitigate these issues, we propose the Softmax-Linked Additive Log-Odds Model (SLALOM), which uses a combination of two scores to quantify the role of input tokens
@@
    We theoretically analyze SLALOM and show that (i) it can be represented by transformers, (ii) it can be uniquely identified from data, and (iii) it is highly efficient to estimate
@@
    Experiments on synthetic and real-world datasets with common language models (LMs) confirm the mismatch between surrogate models and predictive models and underline that single scores like LIME [40] or SHAP [31] do not cover different angles of interpretability, such as human attention or faithfulness metrics, as precisely as scores by SLALOM
@@ 2 Related Work
Explainability for Transformers. Various methods exitst to tackle model explainability [36, 9]. Furthermore, specific approaches have been devised for the transformer architecture [55]: As the attention mechanism at the heart of transformer models is supposed to focus on relevant tokens, it seems a good target for explainability methods. Several works are turning to attention patterns as model explanation techniques. A central attention-based method is put forward by Abnar and Zuidema [1], who propose two methods of aggregating raw attentions across layers, flow and rollout. Brunner et al. [8] focus on effective attentions, which aim to identify the portion of attention weights actually influencing the model’s decision. While these approaches follow a scalar approach considering only attention weights, Kobayashi et al. [28, 29] propose a norm-based vector-valued analysis, arguing that relying solely on attention weights is insufficient and the other components of the model need to be considered. Building on the norm-based approach, Modarressi et al. [34, 35] further follow down the path of decomposing the transformer architecture, presenting global level explanations with the help of rollout. Beyond that, many more attention-based explanation approaches have been put forward [12, 23, 18, 37, 50, 58]. However, it is debatable whether attention scores are suitable for explainabilty [56, 25, 46]. Despite the active research in the area, the recovery property is not considered in these prior works and their scope is often limited to very specific transformer models.
@@
Model-agnostic XAI. Researchers have therefore devised model-agnostic explanations that can be applied without precise knowledge of a model’s architecture. Model-agnostic local explanations like LIME [40], SHAP [31] and others [47, 51, 49, 57, 14, etc.] are a particularly popular class of explanations that are applied to LMs as well [52, 44, 17]. Surrogate models are a common subform [22], which locally approximate a black-box model through a simple, interpretable function.
@@
Linking Models and Explanations. Prior work has distilled the link between classes of surrogate models that can be recovered by explanations [3, 22, Theorem 3]. Notable works in include Garreau and von Luxburg [19] which provides analytical results on different parametrizations of LIME and Bordt and von Luxburg [7] which formalizes the connection between Shapley values and GAMs for classical Shapley values as well as $n$-Shapley values that can also model higher-order interactions. We contribute to the literature by showing that transformers are inherently incapable to represent GAMs or linear models, casting doubts on the applicabililty of LIME and SHAP to transformers. We provide SLALOM, a novel surrogate model that maintains a recovery property for transformers.
@@ 3 Preliminaries
Figure 2: Transformer architecture. In each layer, input embeddings ${\bm{h}}_{i}^{(l-1)}$ for each token $i$ are transformed into output embeddings ${\bm{h}}_{i}^{(l)}$. When detaching the part prior to the classification head (“cls”), we see that the output only depends on the last embedding ${\bm{h}}_{1}^{(L-1)}$ and attention output ${\bm{s}}_{1}$.
@@ 3.1 Input and output representations
In this work, we focus on classification problems of token sequences. For the sake of simplicity, we initially consider a 2-class classification problem with labels $y\in\mathcal{Y}=\{0,1\}$. We will outline how to generalize our approaches to multi-class problems in Section C.1.
The input consists of a sequence of tokens ${\bm{t}}{=}\left[t_{1},\ldots,t_{\lvert{\bm{t}}\rvert}\right]$ where ${\lvert{\bm{t}}\rvert}\in 1,\ldots,C$ is the sequence length that can span at most $C$ tokens (the context length). The tokens stem from a finite size vocabulary $\mathcal{V}$, i.e., $t_{i}\in\mathcal{V},i=1,\ldots,\lvert{\bm{t}}\rvert$. To transform the tokens into a representation amenable to processing with computational methods, the tokens need to be encoded as numerical vectors. To this end, an embedding function ${\bm{e}}:\mathcal{V}\rightarrow\mathbb{R}^{d}$ is used, where $d$ is the embedding dimension. Let ${\bm{e}}_{i}={\bm{e}}(t_{i})$ be the embedding of the $i$-th token such that the entire sentence is embedded in a matrix ${\bm{E}}=[{\bm{e}}_{1},\ldots,{\bm{e}}_{\lvert{\bm{t}}\rvert}]^{\top}\in\mathbb{R}^{{\lvert{\bm{t}}\rvert}\times d}$. The output is given by a logit vector ${\bm{l}}\in\mathbb{R}^{\lvert\mathcal{Y}\rvert}$, such that $\text{softmax}({\bm{l}})$ contains individual class probabilities.
@@ 3.2 The common transformer architecture
Many popular LMs follow the transformer architecture introduced by Vaswani et al. [55] with only minor modifications. We will introduce the most relevant building blocks of the architecture in this section. A complete formalization is given in Section B.1. A schematic overview of the architecture is visualized in Figure 2. Let us denote the input embedding of token $i=1,\ldots,{\lvert{\bm{t}}\rvert}$ in layer $l\in 1,\ldots,L$ by ${\bm{h}}^{(l-1)}_{i}{\in}\mathbb{R}^{d}$, where ${\bm{h}}^{(0)}_{i}{=}{\bm{e}}_{i}$. The core component of the attention architecture is the attention head. Although we only formalize a single head here, our theoretical results cover multiple heads. For each token, a query, key, and a value vector are computed by applying an affine-linear transform to the input embeddings. Keys and queries are projected onto each other and normalized by a row-wise softmax operation resulting in attention weights $\alpha_{ij}\in[0,1]$, denoting how much token $i$ is influenced by token $j$. The attention output for token $i$ can be computed as ${\bm{s}}_{i}=\sum_{j=1}^{{\lvert{\bm{t}}\rvert}}a_{ij}{\bm{v}}_{j}$, where ${\bm{v}}_{j}\in\mathbb{R}^{d_{h}}$ denotes the value vector for token $j$. The final ${\bm{s}}_{i}$ are projected back to dimension $d$ by a projection operator $P:\mathbb{R}^{d_{h}}\rightarrow\mathbb{R}^{d}$ before they are added to the corresponding input embedding ${\bm{h}}_{i}^{(l-1)}$ as mandated by skip-connections. The sum is then transformed by a nonlinear function that we denote by $\text{ffn}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$, finally resulting in a transformed embedding ${\bm{h}}^{(l)}_{i}$. This procedure is repeated iteratively for layers $1,\ldots,L$ such that we finally arrive at output embeddings ${\bm{h}}^{(L)}_{i}$. To perform classification, a classification head $\text{cls}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{\lvert\mathcal{Y}\rvert}$ is put on top of a token at some index $r$ (how this token is chosen depends on the architecture, common choices include $r\in\{1,{\lvert{\bm{t}}\rvert}\}$), such that we get the final logit output ${\bm{l}}=\text{cls}\left({\bm{h}}^{(L)}_{r}\right)$. The logit output is transformed to a probability vector via another softmax operation. Note that in the two-class case, we obtain the log-odds $F({\bm{t}})$ by taking the difference ($\Delta$) between the two logits, i.e., $F({\bm{t}})\coloneqq\log\frac{p(y=1|{\bm{t}})}{p(y=0|{\bm{t}})}=\Delta({\bm{l}})={\bm{l}}_{1}-{\bm{l}}_{0}$.

### v2

@@ Abstract
We address the critical challenge of applying feature attribution methods to the transformer architecture, which dominates current applications in natural language processing and beyond. Traditional attribution methods to explainable AI (XAI) explicitly or implicitly rely on linear or additive surrogate models to quantify the impact of input features on a model’s output. In this work, we formally prove an alarming incompatibility: transformers are structurally incapable of representing linear or additive surrogate models used for feature attribution, undermining the grounding of these conventional explanation methodologies. To address this discrepancy, we introduce the Softmax-Linked Additive Log Odds Model (SLALOM), a novel surrogate model specifically designed to align with the transformer framework. SLALOM demonstrates the capacity to deliver a range of insightful explanations with both synthetic and real-world datasets. We highlight SLALOM’s unique efficiency-quality curve by showing that SLALOM can produce explanations with substantially higher fidelity than competing surrogate models or provide explanations of comparable quality at a fraction of their computational costs. We release code for SLALOM as an open-source project online at https://github.com/tleemann/slalom_explanations.
@@ 1 Introduction
The transformer architecture (Vaswani et al., 2017) has been established as the status quo in modern natural language processing (Devlin et al., 2018; Radford et al., 2018; 2019; Touvron et al., 2023). However, the current and foreseeable adoption of large language models (LLMs) in critical domains such as the judicial system (Chalkidis et al., 2019) and the medical domain (Jeblick et al., 2023) comes with an increased need for transparency and interpretability. Methods to enhance the interpretability of an artificial intelligence (AI) system are developed in the research area of Explainable AI (XAI, Adadi & Berrada, 2018; Gilpin et al., 2018; Molnar, 2019; Burkart & Huber, 2021). A recent meta-study (Rong et al., 2023) shows that XAI has the potential to increase users’ understanding of AI systems and their trust therein. Local feature attribution methods that quantify the contribution of each input to a decision outcome are among the most popular explanation methods, and a variety of approaches have been suggested for the task of computing such attributions (Kasneci & Gottron, 2016; Ribeiro et al., 2016; Sundararajan et al., 2017; Lundberg & Lee, 2017; Covert et al., 2021; Modarressi et al., 2022).
@@
GAM, ensembles, boosting GAM: ✓non-linearities ✗no interactions Removal-based, e.g. Shapley Values, Local GAM approximation
@@
Figure 1: Transformers cannot be well explained through additive models. Left: We exemplarily show the log odds for the outputs of a BERT model and a linear Naïve-Bayes model (“linear”) assigning each word a weight trained on the IMDB movie review dataset. The token colors indicate the weights assigned by the linear model. We pass two sequences to the models independently and in concatenation. For the linear model, the output of the concatenated sequence can be described by the sum, but this is not the case for BERT. We show that this phenomenon is not due to a non-linearity in this particular model but stems from a general incapacity of transformers to represent additive functions. Right: To overcome this difficulty, we propose SLALOM, a novel surrogate model specifically designed to better approximate transformer models.
@@
It remains hard to formally define the contribution of an input feature for non-linear functions. Recent work (Han et al., 2022) has shown that common explanation methods do so by implicitly or explicitly performing a local approximation of the complex black-box function, denoted as $f$, using a simpler surrogate function $g$ from a predefined class $\mathcal{G}$. For instance, Local Interpretable Model-agnostic Explanations (LIME, Ribeiro et al., 2016) or input gradient explanations (Baehrens et al., 2010) use a linear surrogate model to approximate the black-box $f$; the model’s coefficients can be used as the feature contributions.
Surrogate model explanations have the advantage that they directly describe the behavior of the model in the proximity of a specific input, i.e., under small perturbations, a property known as fidelity (Guidotti et al., 2018; Nauta et al., 2023). Fidelity can be quantified through the difference between the prediction model’s outputs and the surrogate model’s outputs (Yeh et al., 2019; Zhou et al., 2019). Explanations that quantitatively describe the prediction model’s output under perturbations with low error have high fidelity.
An implication of models with high representative capacity and high-fidelity explanations is the recovery property: If the true relation $f$ between features and labels in the data is already within the function class $\mathcal{G}$, the model will learn this function and we can effectively reconstruct the original $f$ from the explanations. For example, suppose that the black-box function we consider is of linear form, i.e., $f({\bm{x}})={\bm{w}}^{\top}{\bm{x}}$, and has been correctly learned. In this case, a gradient explanation as well as continuous LIME (C-LIME, Agarwal et al., 2021) will recover the original model’s parameters up to an offset (Han et al., 2022, Theorem 1). Shapley value explanations (Lundberg & Lee, 2017) possess a comparable relationship: It is known that they correspond to the pointwise feature contributions of Generalized Additive Models (GAM, Bordt & von Luxburg, 2023). The significance of recovery properties lies in their role when explanations are leveraged to gain insights into the underlying data. Particularly when XAI is used for scientific applications such as drug discovery (Mak et al., 2023), preserving the path from the input data to the explanation through a learned model is crucial. However, such guarantees can only be provided when surrogate function class $\mathcal{G}$ can effectively mimic the model’s learned relation, at least within some local region.
@@
In this study, we demonstrate that the transformer architecture, the main building block of LLMs such as the GPT models (Radford et al., 2019), is inherently incapable of learning additive models on the input tokens, both theoretically and empirically. By additive models, we refer to models that assign each token a weight. The sum of the individual token weights then gives the output of the additive model. Linear models are a subset of this class. We formally prove that simple encoder-only and decoder-only transformers structurally cannot represent such additive models due to the attention mechanism’s softmax normalization, which necessarily introduces token dependencies over the entire sequence length. An example is illustrated in Figure 1 (left). Our finding that the function spaces represented by additive models and transformers are disjoint when dismissing trivial cases implies that prevalent feature attribution explanations are insufficient to model transformers. They cannot possess high fidelity, i.e., they cannot quantitatively describe these models’ behavior well. This also undermines the recovery property, highlighting a significant oversight in current XAI practices. As our results suggest that the role of tokens cannot be described through a single score, we introduce the Softmax-Linked Additive Log Odds Model (SLALOM, cf. Figure 1, right), which represents the role of each input token in two dimensions: The token value describes the independent effect of a token, whereas the token importance provides a token’s interaction weight when combined with other tokens in a sequence. In summary, our work offers the following contributions beyond the related literature:
@@
    We theoretically and empirically demonstrate that common transformer architectures fail to represent additive and linear models on the input tokens, jeopardizing current attribution methods’ fidelity.
@@
    To mitigate these issues, we propose the Softmax-Linked Additive Log Odds Model (SLALOM), which uses a combination of two scores to quantify the role of input tokens.
@@
    We theoretically analyze SLALOM and show that (i) it can be represented by transformers (i.e., the fidelity property), (ii) it can be uniquely identified from data (i.e., the recovery property), and (iii) it is highly efficient to estimate.
@@
    Experiments on synthetic and real-world datasets with common language models (LMs) confirm the mismatch between surrogate models and predictive models, underline that two scores cover different angles of interpretability, and that SLALOM explanations can be computed that have substantially higher fidelity or efficiency than competing techniques.
@@ 2 Related Work
Explainability for transformers. Various methods exist to tackle model explainability (Molnar, 2019; Burkart & Huber, 2021). Furthermore, specific approaches have been devised for the transformer architecture (Vaswani et al., 2017): As the attention mechanism at the heart of transformer models is supposed to focus on relevant tokens, it seems a good target for explainability methods. Several works turn to attention patterns as model explanation techniques. A central attention-based method is put forward by Abnar & Zuidema (2020), who propose two methods of aggregating raw attentions across layers, flow and rollout. Brunner et al. (2020) focus on effective attentions, which aim to identify the portion of attention weights actually influencing the model’s decision. While these approaches follow a scalar approach considering only attention weights, Kobayashi et al. (2020; 2023) propose a norm-based vector-valued analysis, arguing that relying solely on attention weights is insufficient and the other components of the model need to be considered. Building on the norm-based approach, Modarressi et al. (2022; 2023) further follow down the path of decomposing the transformer architecture, presenting global-level explanations with the help of rollout. Beyond that, many more attention-based explanation approaches have been put forward (Chen et al., 2020; Hao et al., 2021; Ferrando & Costa-jussà, 2021; Qiang et al., 2022; Sun et al., 2023; Yang et al., 2023) and relevance-propagation methods such as LRP have been adapted to the transformer architecture (Achtibat et al., 2024). A drawback with these model-specific explanations remains the implementational overhead that is required to adapt these methods for each architecture. On the formal side, there is no explicit method to quantitatively predict the transformer model’s behavior under perturbations leaving the fidelity of these explanations unclear.
@@
Model-agnostic XAI. In contrast to transformer-specific methods, researchers have devised model-agnostic explanations that can be applied without precise knowledge of a model’s architecture. Model-agnostic local explanations like LIME (Ribeiro et al., 2016), SHAP (Lundberg & Lee, 2017) and others (Shrikumar et al., 2017; Sundararajan et al., 2017; Smilkov et al., 2017; Xu et al., 2020; Covert et al., 2021, etc.) are a particularly popular class of explanations that are applied to LMs as well (Szczepański et al., 2021; Schirmer et al., 2023; Dolk et al., 2022). Surrogate models are a common subform (Han et al., 2022), which locally approximate a black-box model through a simple, interpretable function.
@@
Linking models and explanations. Prior work has distilled the link between classes of surrogate models that can be recovered by explanations (Agarwal et al., 2021; Han et al., 2022, Theorem 3). Notable works include Garreau & von Luxburg (2020), which provides analytical results on different parametrizations of LIME, and Bordt & von Luxburg (2023), which formalizes the connection between Shapley values and GAMs for classical Shapley values as well as $n$-Shapley values that can also model higher-order interactions.
Interpreting feature attributions as linear models. Many feature attribution methods do not use explicit surrogate models, e.g., SHAP or LRP. To predict model behavior under perturbations, they can be intuitively interpreted as a linear surrogate model (cf. Yeh et al., 2019): If feature $i$ has a contribution of $\phi_{i}$, removing $i$ should reduce the model output by $\phi_{i}$ (Achtibat et al., 2024), giving rise to an implicit linear model.
We contribute to the literature by theoretically showing that transformers are inherently incapable to represent additive models, casting doubts on the applicability of local LIME, SHAP, attention-based, and other attribution methods to transformers. These methods fit an explicit additive surrogate model or can be interpreted as one. Our finding that additive scores are insufficient to predict the behavior of transformers leaves us with no method that enables strict fidelity. To bridge this gap, we provide SLALOM, a novel surrogate model with substantially increased fidelity and a recovery property for transformers.
@@ 3 Preliminaries
Figure 2: Transformer architecture. In each layer $l{=}1,\ldots,L$, input embeddings ${\bm{h}}_{i}^{(l-1)}$ for each token $i$ are transformed into output embeddings ${\bm{h}}_{i}^{(l)}$. When detaching the part prior to the classification head (“cls”), we see that the output only depends on the last embedding ${\bm{h}}_{1}^{(L-1)}$ and attention output ${\bm{s}}_{1}$.
@@ 3.1 Input and output representations
In this work, we focus on classification problems of token sequences. For the sake of simplicity, we initially consider a 2-class classification problem with labels $y\in\mathcal{Y}=\{0,1\}$. We will outline how to generalize our approach to multi-class problems in Section C.1.
The input consists of a sequence of tokens ${\bm{t}}{=}\left[t_{1},\ldots,t_{\lvert{\bm{t}}\rvert}\right]$ where ${\lvert{\bm{t}}\rvert}\in 1,\ldots,C$ is the sequence length that can span at most $C$ tokens (the context length). All tokens $t_{i}$ in the sequence ${\bm{t}}$ stem from a finite size vocabulary $\mathcal{V}$, i.e., $t_{i}\in\mathcal{V},\forall i=1,\ldots,\lvert{\bm{t}}\rvert$. To transform the tokens into a representation amenable to processing with computational methods, the tokens need to be encoded as numerical vectors. To this end, an embedding function ${\bm{e}}:\mathcal{V}\rightarrow\mathbb{R}^{d}$ is used, where $d$ is the embedding dimension. Let ${\bm{e}}_{i}={\bm{e}}(t_{i})$ be the embedding of the $i$-th token. The output is given by a logit vector ${\bm{l}}\in\mathbb{R}^{\lvert\mathcal{Y}\rvert}$, such that $\text{softmax}({\bm{l}})$ contains individual class probabilities.
@@ 3.2 The common transformer architecture
Many popular LMs follow the transformer architecture introduced by Vaswani et al. (2017) with only minor modifications. We will introduce the most relevant building blocks of the architecture in this section. A complete formalization is given in Section B.1. A schematic overview of the architecture is visualized in Figure 2. Let us denote the input embedding of token $i=1,\ldots,{\lvert{\bm{t}}\rvert}$ in layer $l\in 1,\ldots,L$ by ${\bm{h}}^{(l-1)}_{i}{\in}\mathbb{R}^{d}$, where ${\bm{h}}^{(0)}_{i}{=}{\bm{e}}_{i}$. The core component of the attention architecture is the attention head. [^1] Although we only formalize a single head here, our theoretical results cover multiple heads as well For each token, a query, key, and a value vector are computed by applying an affine-linear transform to the input embeddings. Keys and queries are projected onto each other and normalized by a row-wise softmax operation resulting in attention weights $\alpha_{ij}\in[0,1]$, denoting how much token $i$ is influenced by token $j$. The attention output for token $i$ can be computed as ${\bm{s}}_{i}=\sum_{j=1}^{{\lvert{\bm{t}}\rvert}}a_{ij}{\bm{v}}_{j}$, where ${\bm{v}}_{j}\in\mathbb{R}^{d_{h}}$ denotes the value vector for token $j$. The final ${\bm{s}}_{i}$ are projected back to dimension $d$ by a projection operator $P:\mathbb{R}^{d_{h}}\rightarrow\mathbb{R}^{d}$ before they are added to the corresponding input embedding ${\bm{h}}_{i}^{(l-1)}$ as mandated by skip-connections. The sum is then transformed by a nonlinear function that we denote by $\text{ffn}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$, finally resulting in a transformed embedding ${\bm{h}}^{(l)}_{i}$. This procedure is repeated iteratively for layers $1,\ldots,L$ such that we finally arrive at output embeddings ${\bm{h}}^{(L)}_{i}$. To perform classification, a classification head $\text{cls}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{\lvert\mathcal{Y}\rvert}$ is put on top of a token at some index $r$ (how this token is chosen depends on the architecture, common choices include $r\in\{1,{\lvert{\bm{t}}\rvert}\}$), such that we get the final logit output ${\bm{l}}=\text{cls}\left({\bm{h}}^{(L)}_{r}\right)$. The logit output is transformed to a probability vector via another softmax operation. Note that in the two-class case, we obtain the log odds $F({\bm{t}})$ by taking the difference ($\Delta$) between the two logits, i.e., $F({\bm{t}})\coloneqq\log\frac{p(y=1|{\bm{t}})}{p(y=0|{\bm{t}})}=\Delta({\bm{l}})={\bm{l}}_{1}-{\bm{l}}_{0}$.